In a cell, the concentration of glucose-6-phosphate is 83 μM and of fructose-6-phosphate is 14 μM. The ΔG°' = 410 cal/mol for the conversion of glucose-6-phosphate to fructose-6-phosphate. Assume the gas constant “R” is 1.987 and T is 298K. Based upon this information, the ΔG’ for this reaction (in cal/mol) is?
Answer #1

For calculating the ΔG’, we can simply use the equation [ΔG°' = ΔG' + RT ln K] where K is the equilibrium constant.

so,

ΔG°' = ΔG' + RT ln K
=> ΔG' = ΔG°' - RT ln K
=> ΔG' = 410 - (1.987*298)[ln(83/14)] = 410-1053.86 = -643.86 cal/mol (Answer)